# ConFIG: Towards Conflict-free Training of Physics Informed Neural Networks

### v1

@@ Abstract
The loss functions of many learning problems contain multiple additive terms that can disagree and yield conflicting update directions. For Physics-Informed Neural Networks (PINNs), loss terms on initial/boundary conditions and physics equations are particularly interesting as they are well-established as highly difficult tasks. To improve learning the challenging multi-objective task posed by PINNs, we propose the ConFIG method, which provides conflict-free updates by ensuring a positive dot product between the final update and each loss-specific gradient. It also maintains consistent optimization rates for all loss terms and dynamically adjusts gradient magnitudes based on conflict levels. We additionally leverage momentum to accelerate optimizations by alternating the back-propagation of different loss terms. The proposed method is evaluated across a range of challenging PINN scenarios, consistently showing superior performance and runtime compared to baseline methods. We also test the proposed method in a classic multi-task benchmark, where the ConFIG method likewise exhibits a highly promising performance. Source code is available at https://tum-pbs.github.io/ConFIG.
@@ 1 Introduction
Efficiently solving partial differential equations (PDEs) is crucial for various scientific fields such as fluid dynamics, electromagnetics, and financial mathematics. However, the nonlinear and high-dimensional PDEs often present significant challenges to the stability and convergence of traditional numerical methods. With the advent of deep learning, there is a strongly growing interest in using this technology to solve PDEs [1, 2]. In this context, Physics Informed Neural Networks (PINNs) [3, 4] leverage networks as a continuous and differentiable ansatz for the underlying physics.
PINNs use auto-differentiation of coordinate-based neural networks, i.e., implicit neural representation (INR), to approximate PDE derivatives. The residuals of the PDEs, along with boundary and initial conditions, are treated as loss terms and penalized during training to achieve a physically plausible solution. Despite their widespread use, training PINNs is a well-recognized challenge [4, 5, 6, 7, 8] due to several possible factors like unbalanced back-propagated gradients from numerical stiffness [6], different convergence rates among loss terms [8], PDE-based soft constraints [7], poor initialization [9], and suboptimal sampling strategies [10]. Traditional methods to improve PINN training typically involve adjusting the weights for PDE residuals and loss terms for initial/boundary conditions [11, 12, 13]. However, although these methods claim to have better solution accuracy, there is currently no consensus on the optimal weighting strategy in PINN training.
Meanwhile, methods manipulating the gradient of each loss term have become popular in Multi-Task Learning (MTL) and Continual Learning (CL) [14, 15, 16] to address conflicts between loss-specific gradients that induce negative transfer [17] and catastrophic forgetting [18]. Unlike weighting strategies that adjust the final update direction solely by modifying each loss-specific gradient’s magnitude, these methods take greater flexibility and often alter the directions of loss-specific gradients. We notice that similar conflicts between gradients also arise in the training of PINNs. First, the gradient magnitude of PDE residuals typically surpasses that of initial/boundary conditions [6], causing the final update gradient to lean heavily towards the residual term. Additionally, PDE residuals often have many local minima due to the infinite number of PDE solutions without prescribed initial/boundary conditions [10]. Consequently, when the optimization process approaches local minima of the PDE residual, the combined update conflicts with the gradient of initial/boundary conditions, severely impeding learning progress.
@@
In the following, we propose the Conflict-Free Inverse Gradients (ConFIG) method to mitigate the conflicts of loss-specific gradients during PINNs training. Our approach provides an update gradient $\bm{g}_{\text{ConFIG}}$ that reduces all loss terms to prevent the optimization from being stuck in the local minima of a specific loss term. With the ConFIG method, the optimization in Fig. 1 converges to the shared minimum for both losses. Our method is characterized by the following properties:
@@ 2 Related Work
Some intuitive weighting strategies come from the penalty method of constrained optimization problems, where higher weights are assigned to less-optimized losses [11, 13]. McClenny et al.  [12] extended this idea by directly setting weights to sample points with soft attention masks. On the contrary, Rafael et al.  [19] advocated equalizing the reduction rates of all loss terms as it guides training toward Pareto optimal solutions from an MTL perspective. Meanwhile, Wang et al.  [6] addressed numerical stiffness issues by determining the weight according to the magnitude of each loss-specific gradient. Wang et al.  [8] explored training procedures via the Neural Tangent Kernel (NTK) perspective, adaptively setting weights based on NTK eigenvalues. Xiang et al.  [20] instead employed Gaussian probabilistic models for uncertainty quantification, estimating weights through maximum likelihood estimation. Others have pursued specialized approaches. E.g., Meer et al.  [21] derived optimal weight parameters for specific problems and devised heuristics for general problems. De Wolff et al.  [22] use an evolutionary multi-objective algorithm to find the trade-offs between individual loss terms in PINNs training. Shin et al.  [23] conducted convergence analyses for PINNs solving linear second-order elliptic and parabolic type PDEs, employing Lipschitz regularized loss. In contrast to adaptive methods, Wight et al.  [24] proposed a fixed weighting strategy for Phase field models. They artificially elevated weights for the loss term associated with initial conditions.
@@
Gradient improvement strategies in MTL and CL focus on understanding and resolving conflicts between loss-specific gradients. Riemer et al.  [14] and Du et al.  [25] proposed using the dot product (cosine similarity) of two gradient vectors to assess whether updates will conflict with each other. Riemer et al. introduced an additional term in the loss to modify the gradient direction, maximizing the dot product of the gradients. Conversely, Du et al. chose to discard gradients of auxiliary tasks if they conflict with the main task. Another intuitive method to resolve conflicts is orthogonal projection. Mehrdad et al.  [15] proposed the Orthogonal Gradient Descent (OGD) method for continual learning, projecting gradients from new tasks onto a subspace orthogonal to the previous task. Chaudhry et al.  [26] proposed learning tasks in low-rank vector subspaces that are kept orthogonal to minimize interference. Yu et al.  [16] introduced the PCGrad method that projects a task’s gradient onto the orthogonal plane of other gradients. Liu et al.  [27] designed the IMTL-G method to ensure that the final gradient has the same projection length on all other vectors. Dong et al.  [28] utilized the Singular Value Decomposition (SVD) on gradient vectors to obtain an orthogonal basis. Javaloy et al.  [29] introduces a method that jointly homogenizes gradient magnitudes and directions for MTL. Few studies in the PINN community have utilized gradient-based improvement strategies. Zhou et al.  [30] introduced the PCGrad method to PINN training for reliability assessment of multi-state systems. Yao et al.  [31] developed a MultiAdam method where they apply Adam optimizer for each loss term separately.
@@ 3.1 Conflict-free Inverse Gradients Method
Generically, we consider an optimization procedure with a set of $m$ individual loss functions, i.e., $\{\mathcal{L}_{1},\mathcal{L}_{2},\cdots,\mathcal{L}_{m}\}$. Let $\{\bm{g}_{1},\bm{g}_{2},\cdots,\bm{g}_{m}\}$ denote the individual gradients corresponding to each of the loss functions. A gradient-descent step with gradient $\bm{g}_{c}$ will conflict with the decrease of $\mathcal{L}_{i}$ if $\bm{g}_{i}^{\top}\bm{g}_{c}$ is negative [14, 25]. Thus, to ensure that all losses are decreasing simultaneously along $\bm{g}_{c}$, all $m$ components of $[\bm{g}_{1},\bm{g}_{2},\cdots,\bm{g}_{m}]^{\top}\bm{g}_{c}$ should be positive. This condition is fulfilled by setting $\bm{g}_{c}=[\bm{g}_{1},\bm{g}_{2},\cdots,\bm{g}_{m}]^{-\top}\bm{w}$, where $\bm{w}=[w_{1},w_{2},\cdots,w_{m}]$ is a vector with $m$ positive components and $M^{-\top}$ is the pseudoinverse of the transposed matrix $M^{\top}$.
@@
We call $\bm{\hat{w}}$ the direction weight. The projection length of $\bm{g}_{c}$ on each loss-specific gradient serves as an effective “learning rate” for each loss. Here, we choose $\hat{w}_{i}=\hat{w}_{j}\ \forall i,j\in[1,m]$ to ensure a uniform decrease rate of all losses, as it was shown to yield a weak form of Pareto optimality for multi-task learning [19].
@@
Here, $\mathbf{1}_{m}$ is a unit vector with $m$ components. A theoretical proof of ConFIG’s convergence in convex and non-convex landscapes can be found in Appendix A.1. The ConFIG method utilizes the pseudoinverse of the gradient matrix to obtain a conflict-free direction. In Appendix A.3, we prove that such an inverse operation is always feasible as long as the dimension of parameter space is larger than the number of losses. Besides, while calculating the pseudoinverse numerically could involve additional computational cost, it is not significant compared to the cost of back-propagation for each loss term. A detailed breakdown of the computational cost can be found in Appendix A.6.
@@
PCGrad [16] and IMTL-G [27] methods from multi-task learning studies also have a similar simplified form for the two-loss scenario. PCGrad projects loss-specific gradients onto the normal plane of others if their cosine similarity is negative, while IMTL-G rescales loss-specific gradients to equalize the final gradient’s projection length on each loss-specific gradient, as illustrated in Fig. 2 and 2. Our ConFIG method in Fig. 2 employs orthogonal components of each loss-specific gradient, akin to PCGrad. Meanwhile, it also ensures the same decrease rate of all losses, similar to IMTL-G. As a result, these three methods share an identical update direction but a different update magnitude in the two-loss scenario. This provides a valuable opportunity to evaluate our adaptive magnitude strategy: for two losses, the PCGrad and IMTL-G methods can be viewed as ConFIG variants with different magnitude rescaling strategies.
@@ 3.3 ConFIG with Momentum Acceleration
Gradient-based methods like PCGrad, IMTL-G, and the proposed ConFIG method require separate backpropagation steps to compute the gradient for each loss term. In contrast, conventional weighting strategies only require a single backpropagation for the total loss. Thus, gradient-based methods are usually $r=\sum_{i}^{m}\mathcal{T}_{b}(\mathcal{L}_{i})/\mathcal{T}_{b}\left(\sum_{i}^{m}\mathcal{L}_{i}\right)$ times more computationally expensive than weighting methods, where $\mathcal{T}_{b}$ is the computational cost of backpropagation. To address this issue, we introduce an accelerated momentum-based variant of ConFIG: M-ConFIG. Our core idea is to leverage the momentum of the gradient for the ConFIG operation and update momentum variables in an alternating fashion to avoid backpropagating all losses in a single step. In each iteration, only a single momentum is updated with its corresponding gradient, while the others are carried over from previous steps. This reduces the computational cost of the M-ConFIG method to $1/m$ of the ConFIG method.
Algorithm 1 details the entire procedure of M-ConFIG. It aligns with the fundamental principles of the Adam algorithm, where the first momentum averages the local gradient, and the second momentum adjusts the magnitude of each parameter. In this approach, we calculate only a single second momentum using an estimated gradient based on the output of the ConFIG operation. An alternative could involve calculating the second momentum for every loss term, similar to the MultiAdam method [31]. However, we found this strategy to be inaccurate and numerically unstable compared to Algorithm 1. A detailed discussion and comparison can be found in the Appendix 1.
@@
Surprisingly, M-ConFIG not only catches up with the training speed of regular weighting strategy in PINNs experiment but typically yields an even lower average computational cost per iteration. This stems from the fact that PINN training often uses a reduced number of sampling points for boundary/initial terms, which are faster to evaluate than the residual term. Thus, $r$ is usually smaller than $m$ for PINNs training. In our experiments, we observed an average value of $r=1.67$ and a speed-up of $1.67/3\approx 0.56$ per iteration for PINNs trained with three losses using M-ConFIG.
@@ 4 Experiments
In this section, we employ the proposed methods for training PINNs on several challenging PDEs. We also explore the application of our method on a classical Multi-Task Learning (MTL) benchmark as an outlook. Unless mentioned otherwise, every result is computed via averaging three training runs initialized with different random seeds, each using the model with the best test performance during the training. For detailed numerical values and standard deviations of the result in each experiment, please refer to the Appendix A.9 and A.10. Training configurations and hyper-parameters of each experiment can be found in the Appendix A.11.

### v2

@@ Abstract
The loss functions of many learning problems contain multiple additive terms that can disagree and yield conflicting update directions. For Physics-Informed Neural Networks (PINNs), loss terms on initial/boundary conditions and physics equations are particularly interesting as they are well-established as highly difficult tasks. To improve learning the challenging multi-objective task posed by PINNs, we propose the ConFIG method, which provides conflict-free updates by ensuring a positive dot product between the final update and each loss-specific gradient. It also maintains consistent optimization rates for all loss terms and dynamically adjusts gradient magnitudes based on conflict levels. We additionally leverage momentum to accelerate optimizations by alternating the back-propagation of different loss terms. We provide a mathematical proof showing the convergence of the ConFIG method, and it is evaluated across a range of challenging PINN scenarios. ConFIG consistently shows superior performance and runtime compared to baseline methods. We also test the proposed method in a classic multi-task benchmark, where the ConFIG method likewise exhibits a highly promising performance. Source code is available at https://tum-pbs.github.io/ConFIG
@@ 1 Introduction
Efficiently solving partial differential equations (PDEs) is crucial for various scientific fields such as fluid dynamics, electromagnetics, and financial mathematics. However, the nonlinear and high-dimensional PDEs often present significant challenges to the stability and convergence of traditional numerical methods. With the advent of deep learning, there is a strongly growing interest in using this technology to solve PDEs (Han et al., 2018; Beck et al., 2023). In this context, Physics Informed Neural Networks (PINNs) (Raissi et al., 2019; Cuomo et al., 2022) leverage networks as a continuous and differentiable ansatz for the underlying physics.
PINNs use auto-differentiation of coordinate-based neural networks, i.e., implicit neural representation (INR), to approximate PDE derivatives. The residuals of the PDEs, along with boundary and initial conditions, are treated as loss terms and penalized during training to achieve a physically plausible solution. Despite their widespread use, training PINNs is a well-recognized challenge (Cuomo et al., 2022; Lino et al., 2023; Wang et al., 2021; Krishnapriyan et al., 2021; Wang et al., 2022) due to several possible factors like unbalanced back-propagated gradients from numerical stiffness (Wang et al., 2021), different convergence rates among loss terms (Wang et al., 2022), PDE-based soft constraints (Krishnapriyan et al., 2021), poor initialization (Wong et al., 2024), and suboptimal sampling strategies (Daw et al., 2023). Traditional methods to improve PINN training typically involve adjusting the weights for PDE residuals and loss terms for initial/boundary conditions (Liu & Wang, 2021; McClenny & Braga-Neto, 2023; Son et al., 2023). However, although these methods claim to have better solution accuracy, there is currently no consensus on the optimal weighting strategy.
Meanwhile, methods manipulating the gradient of each loss term have become popular in Multi-Task Learning (MTL) and Continual Learning (CL) (Riemer et al., 2019; Farajtabar et al., 2020; Yu et al., 2020) to address conflicts between loss-specific gradients that induce negative transfer (Long et al., 2017) and catastrophic forgetting (Kirkpatrick et al., 2017). Unlike weighting strategies that adjust the final update direction solely by modifying each loss-specific gradient’s magnitude, these methods take greater flexibility and often alter the directions of loss-specific gradients. We notice that similar conflicts between gradients also arise in the training of PINNs. First, the gradient magnitude of PDE residuals typically surpasses that of initial/boundary conditions (Wang et al., 2021), causing the final update gradient to lean heavily towards the residual term. Additionally, PDE residuals often have many local minima due to the infinite number of PDE solutions without prescribed initial/boundary conditions (Daw et al., 2023). Consequently, when the optimization process approaches local minima of the PDE residual, the combined update conflicts with the gradient of initial/boundary conditions, severely impeding learning progress.
@@
In the following, we propose the Conflict-Free Inverse Gradients (ConFIG) method to mitigate the conflicts of loss-specific gradients during PINNs training. Our approach provides an update gradient $\bm{g}_{\text{ConFIG}}$ that reduces all loss terms to prevent the optimization from being stuck in the local minima of a specific loss term. With the ConFIG method, the optimization in Fig. 1 converges to the shared minimum for both losses. Our method is provably convergent, and characterized by the following properties:
@@ 2 Related Work
Some intuitive weighting strategies come from the penalty method of constrained optimization problems, where higher weights are assigned to less-optimized losses (Liu & Wang, 2021; Son et al., 2023). McClenny & Braga-Neto (2023) extended this idea by directly setting weights to sample points with soft attention masks. On the contrary, Bischof & Kraus (2021) advocated equalizing the reduction rates of all loss terms as it guides training toward Pareto optimal solutions from an MTL perspective. Meanwhile, Wang et al. (2021) addressed numerical stiffness issues by determining the weight according to the magnitude of each loss-specific gradient. Wang et al. (2022) explored training procedures via the Neural Tangent Kernel (NTK) perspective, adaptively setting weights based on NTK eigenvalues. Xiang et al. (2022) instead employed Gaussian probabilistic models for uncertainty quantification, estimating weights through maximum likelihood estimation. Others have pursued specialized approaches. E.g., van der Meer et al. (2022) derived optimal weight parameters for specific problems and devised heuristics for general problems. de Wolff et al. (2022) use an evolutionary multi-objective algorithm to find the trade-offs between individual loss terms in PINNs training. Shin et al. (2020) conducted convergence analyses for PINNs solving linear second-order elliptic and parabolic type PDEs, employing Lipschitz regularized loss. In contrast to adaptive methods, Wight & Zhao (2021) proposed a fixed weighting strategy for phase field models. They artificially elevated weights for the loss term associated with initial conditions.
@@
Gradient improvement strategies in MTL and CL focus on understanding and resolving conflicts between loss-specific gradients. Riemer et al. (2019) and Du et al. (2018) proposed using the dot product (cosine similarity) of two gradient vectors to assess whether updates will conflict with each other. Riemer et al. (2019) introduced an additional term in the loss to modify the gradient direction, maximizing the dot product of the gradients. Conversely,Du et al. (2018) chose to discard gradients of auxiliary tasks if they conflict with the main task. Another intuitive method to resolve conflicts is orthogonal projection. Farajtabar et al. (2020) proposed the Orthogonal Gradient Descent (OGD) method for continual learning, projecting gradients from new tasks onto a subspace orthogonal to the previous task. Chaudhry et al. (2020) proposed learning tasks in low-rank vector subspaces that are kept orthogonal to minimize interference. Yu et al. (2020) introduced the PCGrad method that projects a task’s gradient onto the orthogonal plane of other gradients. Liu et al. (2021b) designed the IMTL-G method to ensure that the final gradient has the same projection length on all other vectors. Dong et al. (2022) utilized the Singular Value Decomposition (SVD) on gradient vectors to obtain an orthogonal basis. Javaloy & Valera (2022) introduces a method that jointly homogenizes gradient magnitudes and directions for MTL. Quinton & Rey (2024) propose a Jacobian descent method with an aggregation step, which constrains the update gradient into the positive cone in the parameter space to avoid conflicts during training. Few studies in the PINN community have utilized gradient-based improvement strategies. Zhou et al. (2023) introduced the PCGrad method to PINN training for reliability assessment of multi-state systems. Yao et al. (2023) developed a MultiAdam method where they apply Adam optimizer for each loss term separately. Concurrent work from Hwang & Lim (2024) proposed a dual cone gradient descent method to mitigate gradient conflict when training PINNs.
@@ 3.1 Conflict-free Inverse Gradients Method
Generically, we consider an optimization procedure with a set of $m$ individual loss functions, i.e., $\{\mathcal{L}_{1},\mathcal{L}_{2},\cdots,\mathcal{L}_{m}\}$. Let $\{\bm{g}_{1},\bm{g}_{2},\cdots,\bm{g}_{m}\}$ denote the individual gradients corresponding to each of the loss functions. A gradient-descent step with gradient $\bm{g}_{c}$ will conflict with the decrease of $\mathcal{L}_{i}$ if $\bm{g}_{i}^{\top}\bm{g}_{c}$ is negative (Riemer et al., 2019; Du et al., 2018). Thus, to ensure that all losses are decreasing simultaneously along $\bm{g}_{c}$, all $m$ components of $[\bm{g}_{1},\bm{g}_{2},\cdots,\bm{g}_{m}]^{\top}\bm{g}_{c}$ should be positive. This condition is fulfilled by setting $\bm{g}_{c}=[\bm{g}_{1},\bm{g}_{2},\cdots,\bm{g}_{m}]^{-\top}\bm{w}$, where $\bm{w}=[w_{1},w_{2},\cdots,w_{m}]$ is a vector with $m$ positive components and $M^{-\top}$ is the pseudoinverse of the transposed matrix $M^{\top}$.
@@
We call $\bm{\hat{w}}$ the direction weight. The projection length of $\bm{g}_{c}$ on each loss-specific gradient serves as an effective “learning rate” for each loss. Here, we choose $\hat{w}_{i}=\hat{w}_{j}\ \forall i,j\in[1,m]$ to ensure a uniform decrease rate of all losses, as it was shown to yield a weak form of Pareto optimality for multi-task learning (Bischof & Kraus, 2021).
@@
Here, $\mathbf{1}_{m}$ is a unit vector with $m$ components. A mathematical proof of ConFIG’s convergence in convex and non-convex landscapes can be found in Appendix A.1. The ConFIG method utilizes the pseudoinverse of the gradient matrix to obtain a conflict-free direction. In Appendix A.3, we prove that such an inverse operation is always feasible as long as the dimension of parameter space is larger than the number of losses. Besides, while calculating the pseudoinverse numerically could involve additional computational cost, it is not significant compared to the cost of back-propagation for each loss term. A detailed breakdown of the computational cost can be found in Appendix A.6.
@@
PCGrad (Yu et al., 2020) and IMTL-G (Liu et al., 2021b) methods from multi-task learning studies also have a similar simplified form for the two-loss scenario. PCGrad projects loss-specific gradients onto the normal plane of others if their cosine similarity is negative, while IMTL-G rescales loss-specific gradients to equalize the final gradient’s projection length on each loss-specific gradient, as illustrated in Fig. 2 and 2. Our ConFIG method in Fig. 2 employs orthogonal components of each loss-specific gradient, akin to PCGrad. Meanwhile, it also ensures the same decrease rate of all losses, similar to IMTL-G. As a result, these three methods share an identical update direction but a different update magnitude in the two-loss scenario. This provides a valuable opportunity to evaluate our adaptive magnitude strategy: for two losses, the PCGrad and IMTL-G methods can be viewed as ConFIG variants with different magnitude rescaling strategies.
@@ 3.3 ConFIG with Momentum Acceleration
Gradient-based methods like PCGrad, IMTL-G, and the proposed ConFIG method require separate backpropagation steps to compute the gradient for each loss term. In contrast, conventional weighting strategies only require a single backpropagation for the total loss. Thus, gradient-based methods are usually $r=\sum_{i}^{m}\mathcal{T}_{b}(\mathcal{L}_{i})/\mathcal{T}_{b}\left(\sum_{i}^{m}\mathcal{L}_{i}\right)$ times more computationally expensive than weighting methods, where $\mathcal{T}_{b}$ is the computational cost of backpropagation. To address this issue, we introduce an accelerated momentum-based variant of ConFIG: M-ConFIG. Our core idea is to leverage the momentum of the gradient for the ConFIG operation and update momentum in an alternating fashion to avoid backpropagating all losses in a single step. In each iteration, only a single momentum is updated with its corresponding gradient, while the others are carried over from previous steps. This reduces the computational cost of the M-ConFIG method to $1/m$ of the ConFIG method.
Algorithm 1 details the entire procedure of M-ConFIG. It aligns with the fundamental principles of the Adam algorithm, where the first momentum averages the local gradient, and the second momentum adjusts the magnitude of each parameter. In this approach, we calculate only a single second momentum using an estimated gradient based on the output of the ConFIG operation. An alternative could involve calculating the second momentum for every loss term, similar to the MultiAdam method (Yao et al., 2023). However, we found this strategy to be inaccurate and numerically unstable compared to Algorithm 1. A detailed discussion and comparison can be found in the Appendix 1.
@@
Surprisingly, M-ConFIG not only catches up with the training speed of regular weighting strategies but typically yields an even lower average computational cost per iteration. This stems from the fact that backpropagating a sub-loss $\mathcal{L}_{i}$ is usually faster than backpropagating the total loss $\sum_{i}^{m}\mathcal{L}_{i}$. Thus, $r$ is usually smaller than $m$ and $r/m<1$. This is especially obvious for PINN training, where a reduced number of sampling points are used for boundary/initial terms and are faster to evaluate than the residual term. In our experiments, we observed an average value of $r=1.67$ and a speed-up of $1.67/3\approx 0.56$ per iteration for PINNs trained with three losses using M-ConFIG.
@@ 4 Experiments
In this section, we employ the proposed methods for training PINNs on several challenging PDEs. We also explore the application of our method on a classical Multi-Task Learning (MTL) benchmark as an outlook. Unless mentioned otherwise, every result is computed via averaging three training runs initialized with different random seeds, each using the model with the best test performance during the training. For detailed numerical values and standard deviations of the result in each experiment, please refer to the Appendix A.9 and A.10. Training configurations and hyper-parameters of each experiment can be found in the Appendix A.12. An additional ablation study on training hyperparameters can be found in the Appendix A.13. It shows that the gains from ConFIG are robust w.r.t. hyperparameter changes.